# A Theory of Recommendations

### v1

@@ 5.3.2 Prevalence of Controversial Products
Without controversial products, preference heterogeneity would not matter: All receivers would accept recommendations, as they would reveal whether the product is good or bad. Hence, the value of the recommendation system would be independent of the preference distribution. More broadly, we now ask how an increasing prevalence of controversial products (an increase in $Q$) affects the value of a recommendation system. A natural conjecture is that recommendations become less valuable when controversial products are more likely. We find that this conjecture needs to be qualified.
@@
Figure 4 illustrates the result. In the white area in Panel A, $V(\beta)$ is decreasing in $\beta$; in the shaded area, an interior maximizer exists. When good and bad products are similarly likely ($\sigma$ is close to $1$), a higher prevalence of controversial products indeed reduces $V(\beta)$. For very high or low values of $\sigma$, however, this need not be true. This surprising result reflects the multiplicative structure of $V(\beta)$. Under Assumption 1, $V(\beta)$ is the product of the buying probability $\pi^{B}$ and the objective value of the recommendation per buyer, $\Delta_{O}^{B}$. The effect of $Q$ on $\pi^{B}$ is negative only if $\sigma$ is sufficiently large (see Panel B). Intuitively, as $Q$ increases, good products are less likely, reducing the probability of buy recommendations. However, there are now more buy recommendations stemming from controversial products, as these have become more likely. $\beta$ and $\sigma$ determine which of these effects dominates. When $\beta$ is high, the increase induced by controversial products leads to more buy recommendations, strengthening the positive effect. When $\sigma$ is high, the lower share of good products dominates, strengthening the negative effect. Similarly, one can show that the effect of $Q$ on $\Delta_{O}^{B}$ becomes negative only for small $\sigma$ (see Panel C). Again, this reflects countervailing forces. The negative effect results from the increased probability that a product with a buy recommendation is controversial. The positive effect reflects a change in the value of the outside option. The values of $\sigma$ and $\beta$ determine which effect dominates. [^28] When $\beta$ is small, the negative effect is weak, as only few types give buy recommendations for controversial products. When $\sigma$ is small, the outside option becomes more valuable, as bad products become less likely, weakening the positive effect.
@@ 6.1 Distinct Sender and Receiver Distributions
We modify the benchmark model by assuming that the sender and the receiver are drawn from different distributions $F$ and $G$, respectively. For instance, this enables us to take into account that the “early adopters” of a new product may have different preferences than the general population (similar to Acemoglu et al. (2022)). We show that the results of the benchmark case easily generalize to this setting.
@@
Consider the model with distinct preference distributions of senders and receivers. Suppose the sender distribution $F$ satisfies Assumption 1. Then, all types accept the recommendation. Further, $V(R)=\pi^{B}_{F}[\Delta_{O,F}^{B}-\Delta_{S,F}^{B}\mathbb{E}_{G}[i]]$ is
@@
Recall that in the benchmark model symmetry of $F$ implies that all receivers follow the recommendation and the subjective effect of the recommendation cancels out. This is still true if $F$ is distinct from $G$ as long as $G$ is also symmetric and hence $\mathbb{E}_{G}[i]=0$. For more general distributions $G$, all receivers still accept the recommendation for the same reason as above. However, as the receiver distribution is not symmetric, the subjective effect does not cancel out anymore. Even so, the result qualitatively mirrors Proposition 5, as changing the threshold still leads to the same trade-off: Reducing the threshold increases the probability of a buy recommendation but decreases its value conditional on that recommendation. However, in the latter effect, the value is no longer purely objective but consists of the objective part $\Delta_{O}^{B}$ and a subjective part given by $-\Delta_{S}^{B}\mathbb{E}_{G}[i]$, reflecting the receiver distribution $G$. Further, as the subjective effect now influences the value of the system, so does the prevalence of controversial products by affecting the tipping point at which the designer opts for a high or a low threshold. In any case, when $\mathbb{E}_{G}[i]=0$, we obtain the result from the baseline model as a special case. Similarly, the other results in the benchmark model remain (qualitatively) valid. With asymmetrically distributed populations, Propositions 6 and 7 hold for arbitrary receiver distributions as long as we impose Assumption 2 on the sender distribution.
@@ 6.2 Multiple Recommendation Thresholds
We now analyze the effects of reducing the coarseness of the recommendation system. We maintain all assumptions of the benchmark case, except that there are two threshold levels $R_{1},R_{2}\in(0,1)$ with $R_{2}\geq R_{1}$ and three types of recommendations $r\in\{D,N,B\}$ which we interpret as “don’t-buy,” “neither buy nor don’t-buy,” and “buy.” The sender gives a buy recommendation if the payoff is at least $R_{2}$, a don’t-buy recommendation if it is below $R_{1}$, and neither if it is in $[R_{1},R_{2})$. We maintain Assumption 1.
@@
We say that a receiver accepts a recommendation if he buys if $r=B$ and never buys if $r=D$, allowing for either behavior if $r=N$. Assumption 1 again implies that the receiver always buys a product with $r=B$ and never buys if $r=D$. As noted above, $r=N$ rules out both uncontroversial products. Given the symmetric population, $r=N$ leaves the relative likelihood of the controversial products unchanged. Hence, whether a type $i$ wants to buy a product with $r=N$ depends on (i) how likely a good product is and (ii) how much he (dis-)likes the more likely controversial product. As in the benchmark model, we can derive the indifferent type so that, depending on parameters, either all, only sufficiently aligned, or no types buy the product with an intermediate recommendation.
$V(R)$ now consists of two components. First, as in the benchmark, all types accept $r=B$, resulting in the component $\pi^{B}\Delta_{O}^{B}$. [^30] As before, following don’t-buy recommendations does not create value. Moreover, $r=N$ can create value for the types that buy following that neutral recommendation. As noted above, parameter values determine which types do this. Despite this ambiguity, we can formulate the following result.
@@ 7 Conclusion
We have set out to study the role of preference heterogeneity in recommendations, asking questions such as: Under what circumstances do people follow a recommendation? What determines the value of a recommendation system and when is it maximized? Our analysis highlights the importance of disentangling objective information from subjective preferences. We have seen that, depending on the preference distribution, the subjective content can reinforce the positive news of a buy recommendation or reverse it so that a buy recommendation is treated as a bad signal. Conversely, the subjective content can turn a don’t-buy recommendation into good news for some people. The design of a recommendation system, as summarized in the threshold $R$, should thus take into account how it influences the relative objectivity of recommendations.
@@ A.1 Proofs and Technical Details for the Benchmark Model
We shall show that conditions (16) and (17) are equivalent by proving the following two steps.

### v2

@@ 5.3.2 Prevalence of Controversial Products
Without controversial products, preference heterogeneity would not matter: All receivers would accept all recommendations, as they would reveal whether the product is good or bad. Hence, the value of the recommendation system would be independent of the preference distribution. More broadly, we now ask how an increasing prevalence of controversial products (an increase in $Q$) affects the value of a recommendation system. A natural conjecture is that recommendations become less valuable when controversial products are more likely. We find that this conjecture needs to be qualified.
@@
Figure 4 illustrates the result. In the white area in Panel A, $V(\beta)$ is decreasing in $\beta$; in the shaded area, an interior maximizer exists. When good and bad products are similarly likely ($\sigma$ is close to $1$), a higher prevalence of controversial products indeed reduces $V(\beta)$. For very high or low values of $\sigma$, however, this need not be true. This surprising result reflects the multiplicative structure of $V(\beta)$. Under Assumption 1, $V(\beta)$ is the product of the buying probability $\pi^{B}$ and the objective value of the recommendation per buyer, $\Delta_{O}^{B}$. The effect of $Q$ on $\pi^{B}$ is negative only if $\sigma$ is sufficiently large (see Panel B). Intuitively, as $Q$ increases, good products are less likely, reducing the probability of buy recommendations. However, there are now more buy recommendations stemming from controversial products, as these have become more likely. $\beta$ and $\sigma$ determine which of these effects dominates. When $\beta$ is high, the greater likelihood that controversial products receive buy recommendations increases the total probability of buy recommendations, strengthening the positive effect. When $\sigma$ is high, the lower share of good products dominates, strengthening the negative effect. Similarly, one can show that the effect of $Q$ on $\Delta_{O}^{B}$ becomes negative only for small $\sigma$ (see Panel C). Again, this reflects countervailing forces. The negative effect results from the increased probability that a product with a buy recommendation is controversial. The positive effect reflects a change in the value of the outside option. The values of $\sigma$ and $\beta$ determine which effect dominates. [^31] When $\beta$ is small, the negative effect is weak, as only few types give buy recommendations for controversial products. When $\sigma$ is small, the outside option becomes more valuable, as bad products become less likely, weakening the positive effect.
@@ 6.1 Distinct Sender and Receiver Distributions
We modify the benchmark model by assuming that the sender and the receiver are drawn from different distributions $F$ and $G$, respectively. For instance, this enables us to take into account that the “early adopters” of a new product may have different preferences than the general population (similar to Acemoglu et al. (2022)). We show that the results of the benchmark case easily generalize to this setting. More importantly, the general setting allows us to treat the choice of the sender distribution as a potentially interesting design variable.
@@
Consider the model with distinct preference distributions of senders and receivers. Suppose the sender distribution $F$ satisfies Assumption 1. Then, all types accept all recommendations. Further, $V(R)=\pi^{B}_{F}[\Delta_{O,F}^{B}-\Delta_{S,F}^{B}\mathbb{E}_{G}[i]]$ is
@@
Recall that in the benchmark model symmetry of $F$ implies that all receivers accept all recommendations and the subjective effect of the recommendation cancels out. This is still true if $F$ is distinct from $G$ as long as $G$ is also symmetric and hence $\mathbb{E}_{G}[i]=0$. For more general distributions $G$, all receivers still accept all recommendations for the same reason as above. However, as the receiver distribution is not symmetric, the subjective effect does not cancel out anymore. Even so, the result qualitatively mirrors Proposition 5, as changing the threshold still leads to the same trade-off: Reducing the threshold increases the probability of a buy recommendation but decreases its value conditional on that recommendation. However, in the latter effect, the value is no longer purely objective but consists of the objective part $\Delta_{O}^{B}$ and a subjective part given by $-\Delta_{S}^{B}\mathbb{E}_{G}[i]$, reflecting the receiver distribution $G$. Further, as the subjective effect now influences the value of the system, so does the prevalence of controversial products by affecting the tipping point at which the designer opts for a high or a low threshold. In any case, when $\mathbb{E}_{G}[i]=0$, we obtain the result from the baseline model as a special case. Similarly, the other results in the benchmark model remain (qualitatively) valid. With asymmetrically distributed populations, Propositions 6 and 7 hold for arbitrary receiver distributions as long as we impose Assumption 2 on the sender distribution.
Allowing for differences in sender and receiver distributions makes it possible to address another important design issue. In some cases, it is natural to assume that a designer, for instance a platform, can influence the sender distribution or at least the receiver’s perception of it. Assuming that the platform has better knowledge about sender types, it may choose to share some of this information with the receivers or even restrict the option of sending recommendations to certain types of senders (and inform the receivers accordingly). For instance, many platforms such as Amazon inform browsing customers that “customers who bought this item also bought” another item.
There are different ways to model this. First, suppose that $F$ is symmetric and that the designer can render the sender distribution more or less precise using mean-preserving contractions or mean-preserving spreads, respectively. Intuitively, one would expect more precise knowledge about the sender to increase the value of the recommendation system. However, the logic of Proposition 9 shows that this is not necessarily the case. As in Section 5.3.1, the effects of mean-preserving contractions and spreads can be captured as changes in $\beta=F(1/2-R)$, just like changes in $R$. With mild adaptions, the proof of Proposition 9 can thus be used to show that whether more or less precise information about the sender increases the value of recommendations depends on the parameters.
Second, suppose that the sender and receiver distributions are fixed but that the designer can provide the receivers with a signal about the sender that allows for inference about the sender’s type. To make the point as starkly as possible, suppose the signal fully reveals the sender’s type. If the sender is revealed to be type $i=0$, all types accept all recommendations, as both types of recommendations contain only objective information. In contrast to Proposition 2—where this behavior is shown to ensue with extreme thresholds—this would arise for all thresholds $R\in(0,1)$. Intuitively, as the type $i=0$ is indifferent between controversial products, her recommendation never reveals subjective information. In contrast, if the signal reveals a sender of type $i\neq 0$, the receiver will be able to learn about controversial products, too. [^33] For instance, if $i=1$ and $R\in(0,1)$, a buy recommendation reveals that the product is either of version $(1,1)$ or $(1,0)$ Thus, having a sender who is partial to some controversial product is not necessarily a bad thing.
Finally, suppose that the sender and receiver distributions are not independent but somehow correlated. Recommendations would then be “tailored” to the receiver $\bar{i}$, as they would be drawn from a sender distribution $F(i|\bar{i})$, resulting in type-specific posteriors $p^{r}_{j}(i)$ for $j\in\{H,1,2,L\}$. In such a setting, an analogous result to Proposition 1 still holds, establishing that either all all or no recommendations are accepted. Further analysis would give rise to similar complications as in Section 5.2.2 with asymmetric distributions, as the the correlation between sender and receiver distribution would affect the set of types who accept all recommendations.
@@ 6.2 Multiple Recommendation Thresholds
We now analyze the effects of reducing the coarseness of the recommendation system. We maintain all assumptions of the benchmark case, except that there are two threshold levels $R_{1},R_{2}\in(0,1)$ with $R_{2}\geq R_{1}$ and three types of recommendations $r\in\{D,N,B\}$ which we interpret as “don’t-buy,” “neither buy nor don’t-buy,” and “buy.” [^34] Such a two-threshold structure is found to be optimal in models of strategic communication in the context of social learning by Smirnov and Starkov (2024) and Bénabou and Vellodi (2024). The sender gives a buy recommendation if the payoff is at least $R_{2}$, a don’t-buy recommendation if it is below $R_{1}$, and neither if it is in $[R_{1},R_{2})$. We maintain Assumption 1.
@@
We say that a receiver accepts all recommendations if he buys if $r=B$ and does not buy if $r=D$, allowing for either behavior if $r=N$. Assumption 1 again implies that the receiver always buys a product with $r=B$ and never buys if $r=D$. As noted above, $r=N$ rules out both uncontroversial products. Given the symmetric population, $r=N$ leaves the relative likelihood of the controversial products unchanged. Hence, whether a type $i$ wants to buy a product with $r=N$ depends on (i) how likely a good product is and (ii) how much he (dis-)likes the more likely controversial product. As in the benchmark model, we can derive the indifferent type so that, depending on parameters, either all, only sufficiently aligned, or no types buy the product with an intermediate recommendation.
$V(R)$ now consists of two components. First, as in the benchmark, all types accept $r=B$, resulting in the component $\pi^{B}\Delta_{O}^{B}$. [^35] As before, accepting don’t-buy recommendations does not create value. Moreover, $r=N$ can create value for the types that buy following that neutral recommendation. As noted above, parameter values determine which types do this. Despite this ambiguity, we can formulate the following result.
@@ 7 Conclusion
We have set out to study the role of preference heterogeneity in recommendations, asking questions such as: Under what circumstances do people accept a recommendation? What determines the value of a recommendation system and when is it maximized? Our analysis highlights the importance of disentangling objective information from subjective preferences. We have seen that, depending on the preference distribution, the subjective content can reinforce the positive news of a buy recommendation or reverse it so that a buy recommendation is treated as a bad signal. Conversely, the subjective content can turn a don’t-buy recommendation into good news for some people. The design of a recommendation system, as summarized in the threshold $R$, should thus take into account how it influences the relative objectivity of recommendations.
@@ A.1 Proofs and Technical Details for the Benchmark Model
We shall show that conditions (10) and (11) are equivalent, thus proving the claim that receivers either accept all recommendations or never accept recommendations. To do so, we prove the following two steps.